OpenStudy (brekathleen):

HELP ME PLEASE ILL GIVE YOU ANYTHING AHHhhhH ASAP Give the function f(x)= 2(3)^x , Section A is from x=0 to x=1 and Section B is from x=2 to x=3 Part A . Find the average rate of change of each question. Part B. How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. PLEASEEEEE

Nnesha (nnesha):

*ANYTHING* ?

Nnesha (nnesha):

i want CHOCOLATES.

Nnesha (nnesha):

to find the slope you need two points for section A substitute x for 0 and then for 1 to find y values when x =0 y = ?? when x =1 y= ??

Nnesha (nnesha):

same with section b find the y value by given x=2 x=3

Nnesha (nnesha):

like for first one \[\rm f(0)= 2(3)^0\] solve right side f(0)= ?? remember anything to the zero power equal to one

OpenStudy (brekathleen):

@nnesha ahhhhh this is so confusing.

Nnesha (nnesha):

what you don't understand ??

Nnesha (nnesha):

f(x) is same as y so the equation can be written as y =2(3)^x now just substitute x for 0 and for 1 to solve for y